Convert normal blood pressure readings of 120 over 80 mm Hg to newtons per meter squared using the relationship for pressure due to the weight of a fluid (P=phg) rarher than a conversion factor
1
Expert's answer
2019-09-17T12:04:55-0400

The deinsity of mercury is


"\rho=13600\frac{kg}{m^3}"

"P_1=\rho g h_1=(13600)(9.8)(0.12)=16000\frac{N}{m^2}"

"P_2=\rho g h_2=(13600)(9.8)(0.08)=10700\frac{N}{m^2}"

The normal blood pressure lies between these two values.